OpenStudy (anonymous):

find the value of 3x/2+1/x=3x/4

OpenStudy (anonymous):

here

OpenStudy (anonymous):

the answer will be so on : dear briiiibabyxy3

OpenStudy (anonymous):

we will first solve LHS

OpenStudy (anonymous):

LCM IS 2x

OpenStudy (anonymous):

like i told you..

OpenStudy (anonymous):

yes it wil be now become

OpenStudy (anonymous):

(3x^2 + 2)/2x

OpenStudy (anonymous):

now transpose 2x to RHS

OpenStudy (anonymous):

we will get 3x/4 * 2x 6x^2/4

OpenStudy (anonymous):

you cant just bring thte 2x over by itself

OpenStudy (anonymous):

we get 3x^2 + 2 = 6x^2/4

OpenStudy (anonymous):

now transpose 4 to LHS

OpenStudy (anonymous):

we will get

OpenStudy (anonymous):

12x^2 + 8

OpenStudy (anonymous):

now 12x^2 + 8 = 6x^2/4

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

12x^2 + 8 = 6x^2 now subtracting both sides by 6x^2 we will get 12x^2 - 6x^2 + 8

OpenStudy (anonymous):

you can post your answer i need to go to sleep it is 546 a.m

OpenStudy (anonymous):

now subtracting both sides by 8

OpenStudy (anonymous):

its being overed wait

OpenStudy (anonymous):

we will get 6x^2 = -8

OpenStudy (anonymous):

x^2 = -8/6

OpenStudy (anonymous):

-8/6 = -4/3

OpenStudy (anonymous):

x = square root of (-4/3)

OpenStudy (anonymous):

i live in INDIA